OpenStudy (anonymous):

Simplify (4-√3)^2 A. 19 - 8√3 B. 19 + 8√3 C. 16 + 8√3 D. 5 - 4√3 Can anyone help?

OpenStudy (jdoe0001):

\(\bf (4-\sqrt{3})^2\implies (4-\sqrt{3})(4-\sqrt{3})\impliedby FOIL\ it\)

OpenStudy (anonymous):

Ohh, okay. I see. Thanks. So, it'd be A?